The main task for me on this maintenance trip was to replace the Dec worm gear and attempt to resolve the homing issues that we have been having in the Dec axis.
|New (right) and old (left) Dec worms for comparison|
At this point we found that a certain member of our team (who shall remain nameless) had been a little overly generous greasing the Dec worm... as a result the Dec homing sensor had been blocked with grease.
Stripping the drive again to clean the homing sensor was easy enough, but time consuming.
Once again we had the whole system assembled and well balanced.
When we tried to home the mount to check the clearances around the scope we found that we had acquired a worrying vibration in the Dec axis drive when the system was slewing at speed. It was not vibrating at tracking speeds, but whilst homing and slewing around the sky the vibration was quite alarming.
At this point we had two frustrating days of the Observatory Open Doors event where we could not continue to work on the system.
After the Open Doors event and with help from the mount manufacturer, Software Bisque, we found the problem and remedied the vibration. As an added bonus we now have our Dec motor drive belts very well tensioned as well... this wasn't the problem, but it was one of the steps we took in the diagnosis.
After a night of alignment work, the telescope returned to operation on the night of the 25/26th. Sadly the result of our efforts to remedy the homing issue appear to have been negligible.
We spent a long whilst yesterday working to understand the nature of the error. We are now certain that the error is related to the mount, and definitely not a problem of slop or flex in our imaging system. We are currently waiting for some more hints from Software Bisque as we are now at an impasse.
The system will continue to operate with the plate solving system in place rejecting images with poor pointing.